What line resembles y=0.5*x+2
Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52781)

.
Straight line.

Did you think differently ?


Answer by addingup(3677)

set x to 0:
y = 0.5(0)+2 any number times 0 equals 0
y = 0+2
y = 2
this gives us our first ordered pair (x, y): (0, 2)
Now set y to 0
-0.5x+y = 2
-0.5x+0 = 2 any number +0 = the number
-0.5x = 2
x = -4
So we have our second ordered pair (x, y): (-4, 0)
Graph these two points and you have your line.